USING HOT EMISSION SPECTRA IN GENERATING LINE LISTS OF MOLECULES (NH_{3}, CH_{4}) FOR ASTROPHYSICAL APPLICATIONS

R.J. HARGREAVES, L. MICHAUX, Department of Chemistry, University of York, Heslington, York, YO10 5DD, UK; G. LI, Harvard-Smithsonian Center for Astrophysics, Atomic and Molecular Physics Division, MS#50, 60 Garden St., Cambridge, MA, 02138,USA; C. BEALE, Department of Chemistry & Biochemistry, Old Dominion University, 4541 Hampton Boulevard, Norfolk, VA, 23529-0126, USA; M. IRFAN, School of Physics and Astronomy, The University of Manchester, Oxford Road, Manchester, M13 9PL, UK; and \underline{P.F. BERNATH}, Department of Chemistry & Biochemistry, Old Dominion University, 4541 Hampton Boulevard, Norfolk, VA, 23529-0126, USA; and University of York, Dept. of Chemistry, Heslington, York, YO10 5DD, UK.

Spectra of cool stars, brown dwarfs and extrasolar planets (exoplanets) contain a dense forest of lines from hot molecules. Examples include CH_{4} and NH_{3} in brown dwarfs and CH_{4} in ‘hot Jupiter’ exoplanets. These observations present challenges to astronomers, who typically use databases such as HITRAN intended for room-temperature applications, to model the spectral energy distributions. We have used a novel technique to combine ‘hot’ emission spectra recorded for a range of sample temperatures (300 – 1400^{∘}C) in order to deduce empirical lower state energies of the emitted lines. We have applied this method to NH_{3} in the 740 – 2100 cm^{−1} range^{a} which includes the ν_{2} and the ν_{4} fundamental modes and in the 1650 – 4000 cm^{−1} range^{b} which includes the ν_{1} and ν_{3} fundamental modes. We have estimated empirical lower state energies and our values have been incorporated into the line lists along with line positions and calibrated line intensities. This method is currently being extended to CH_{4}. Our results can be used directly for the simulation of astronomical spectra.

aHargreaves, R. J., Li, G., and Bernath, P. F. Astrophys. J. 735 (2011) 111.

bHargreaves, R. J., Li, G., and Bernath, P. F. J. Quant. Spectrosc. Radiat. Transfer, (2012) in press.